OpenStudy (anonymous):

Solve each radical equation. Check your solution. Square root 5x^2-3x=2x

OpenStudy (***[isuru]***):

is it \[\sqrt{5x ^{2} - 3x } = 2x \] or \[\sqrt{5x ^{2}} -3x = 2x\] or \[\sqrt{5}x ^{2} - 3x = 2x\] I think it's the first one ?

OpenStudy (anonymous):

How do you solve it and yes it is the first one and how do you solve it

OpenStudy (anonymous):

\[\sqrt{5x^2-3x}=2x\]

OpenStudy (anonymous):

Here how to do it:

OpenStudy (anonymous):

First square root each side to remove the square root, which means you gonna have to use the square method on "2x" which equals "4x^2"